The midpoint of sides of a triangle are (-3,2)(1,1) and (5,7).then how find the vertices

Grade:12

1 Answers

Arun
5 years ago
Dear student
 
Let the vertices are (x1 , y1) , (x2 , y2) and (x3 , y3)
 
Then
 
x1 + x2 / 2 = – 3
y1 + y2 /2 = 2
 
x2 + x3 /2 = 1
y2 + y3 /2 = 1
 
x1 + x3/2 = 5
 
y1 + y3 /2 = 7
 
Now you can solve them